# Bending parameterization of one-sided degenerated Kleinian surface groups

### v1

@@ 1. Introduction
Refer to caption$\partial_{\infty}^{+}M$$\partial_{\infty}^{-}M$$C(M)$$\cong S$$C(M)$$\partial^{+}C(M)$$\partial^{-}C(M)$cuspdegenerate end
Figure 1. Schematic representation of hyperbolic manifolds homeomorphic to $S\times{\mathbb{R}}$, with $1$-dimensional “$S$-direction”, conformal boundary in red and convex core in grey. The structures are, from left to right: Fuchsian, quasi-Fuchsian, Kleinian with a cusp and Kleinian with simply-degenerated bottom end.
@@
More precisely, let $P\subset S^{-}$ be a simple multicurve, let $\Sigma^{-}$ be a union of connected components of $S^{-}-P$ and let $L$ be a union of arational laminations on each components of $S^{-}-(P\cup\Sigma^{-})$. Denote by ${\mathcal{AH}}_{o}^{+}(S;P\cup L)\subset{\mathcal{AH}}_{o}(S)$ the subspace of one-sided degenerated structures with bottom end cusped along $P$ and having ending laminations $L$. Let <${\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})\subset{\mathcal{ML}}(S^{+})$ be the subspace of measured laminations $\mu$ on $S^{+}$ without closed leaves of weight $\geq\pi$ and such that $\mu$ and $P\cup L$ fill $S-\Sigma^{-}$, i.e. any essential closed curve on $S$ intersects either $\mu\cup P\cup L$ or $\Sigma^{-}$.
@@ 1.1. Outline of the proof
First, we prove continuity in Section 3, see Proposition 3.4. Then, in section 4 we prove properness. The argument follows the same outline as for similar maps, considered by Bonahon-Otal [BO04], Lecuire [Lec06], Anderson-Lecuire [AL13] and Baba-Oshika [BO23]. We first show that any sequence of representations with convergent image under $\operatorname{m}^{-}\times\operatorname{b}^{+}$ must have an algebraically convergent subsequence, using the theory of degenerescence of metrics towards actions on real trees of Morgan-Shalen [MS84, Ota96]. Then, it follows from results of Anderson-Lecuire [AL13] and Bonahon-Otal and Lecuire [BO04, Lec06] that this algebraic limit must lie in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
@@
[…]
${{[0,\infty)\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{QF}}_{o}(S)}}$${{{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})}}$${{[0,\infty]\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{AH}}_{o}^{+}(S)}}$${{{\mathcal{ML}}(S^{+})}}$$\scriptstyle{r\times\rm{incl}}$$\scriptstyle{\operatorname{U}}$$\scriptstyle{\operatorname{b}^{+}\times\operatorname{b}^{-}}$$\scriptstyle{\rm{proj}_{2}}$$\scriptstyle{\overline{r}\times\rm{incl}}$$\scriptstyle{\operatorname{B}^{+}}$¯¯$\scriptstyle{\overline{\operatorname{U}}}$$\scriptstyle{\operatorname{b}^{+}}$
@@ 1.2. Further directions
The essential difficulty lies in proving that any geometrically finite structure on $S\times{\mathbb{R}}$ is uniquely determined by its pair of bending laminations. This is expected to hold, and it was proved in the case of rational bending laminations by Bonahon and Otal [BO04]. But it is not yet known in the general case, where our strategy fails.
@@ 2.2. Laminations and trees
A transverse measure on a geodesic lamination $L\in{\mathcal{GL}}(S)$ is a map $\mu$ associating to each closed arc transverse to $L$ a number $\mu(k)\in{\mathbb{R}}_{\geq 0}$ (also denoted $i(\mu,k)$) which is invariant under transverse isotopy of arcs, countably additive, and with full support, i.e. $\mu(k)>0$ if and only if $k\cap L\not=\emptyset$. Let ${\mathcal{ML}}(S)$ denote the space of measured lamination on $S$, endowed with the weak topology. The geodesic laminations that admit a transverse measure (with full support) are precisely the recurrent laminations [PH92, §1.7]. A minimal geodesic lamination admits a unique transverse measure up to scaling. See [PH92] for more details on measured laminations, and a description of a piecewise linear structure on ${\mathcal{ML}}(S)$ using train tracks. By Thurston earthquake theorem [Thu86], the space ${\mathcal{ML}}(S)$ is homeomorphic to ${\mathcal{T}}(S)$. See also [CEM06], [Ota96, Appendix], [MT98, §6.1].
@@ 2.4. Pleated surfaces
Let $M$ be a hyperbolic three-manifold, $S$ a complete hyperbolic surface (not necessarily compact) and $L$ a geodesic lamination on $S$. A pleated surface with pleating locus $L$ is a continuous map $f\colon S\to M$ which is a totally geodesic immersion on $S-L$ and on $L$. In general, the flat pieces of $S-L$ are bent along the pleating locus, with an angle recorded into a bending cocycle $\beta$ which is an element of the space ${\mathcal{H}}(L;{\mathbb{R}}/2\pi{\mathbb{Z}})$ of ${\mathbb{R}}/2\pi{\mathbb{Z}}$-valued cocycles along $L$ [Bon96] (which is a torus whose dimension is determined by $S$ and $L$). When $f$ is locally convex, this bending cocycle defines a transverse measure with support contained in $L$, called the bending or pleating (measured) lamination of $f$. The induced path metric on $S$ inside $M$ is hyperbolic and defines a point $h\in{\mathcal{T}}(S)$, which can also be encoded in a shearing cocycle $\sigma\in{\mathcal{H}}(L;{\mathbb{R}})$. The boundary of the convex core $C(\rho)$ of $\rho\in{\mathcal{AH}}(S)$ is an embedded convex pleated surface [CEM06, II.1.11,12]. In particular, its bending measured lamination has weight $<\pi$ along closed leaves.
@@ 2.7. End structures
Baba and Ohshika described the combination of ending laminations, parabolic locus and bending laminations that occur for Kleinian surface groups [BO23]. It follows from their result that, for $\rho\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ where $e^{-}=P\cup L$, the bending lamination $\operatorname{b}^{+}(\rho)$ of the top boundary of its convex core fills $S(e^{-})$ with $e^{-}$, i.e. any essential closed curve on $S(e^{-})$ intersects $P\cup L$ or $\operatorname{b}^{+}(\rho)$ (in other words, any essential closed curve on $S$ intersects $P\cup L\cup\Sigma^{-}\cup\operatorname{b}^{+}(\rho)$). Hence it lies in the subspace <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ consisting of those measured laminations without closed leaves of weight $\geq\pi$, and which fill $S(e^{-})$ with $e^{-}$.
@@ 2.8. Metrics converging to an end structure
[…]
@@
[…]
@@ 4. Properness
The goal of this section is to prove that the map $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper (see the end of Section 2.7 for a description of the bending laminations that occur for strucutes in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$). This will imply that $\operatorname{b}^{+}_{\infty}\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper too. The proof is similar to the proof of properness of similar maps by Baba and Ohshika in [BO23] and Anderson and Lecuire in [AL13].
@@
Let <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ be the subspace of ${\mathcal{ML}}(S^{+})$ consisting of the measured laminations $\mu$ satisfying the following conditions :
@@ 4.2. Algebraic convergence
Write $(\operatorname{m}^{-}\times\operatorname{b}^{+})(\rho_{n})=(m_{n}^{-},\lambda^{+}_{n})$ for each $n\in{\mathbb{N}}$, and $(m^{-},\lambda^{+})$ their limit.
@@
Finally, Bridgeman proved that there is a constant bounding the length of bending laminations of convex cores [Bri98, Prop. 2] (see §2.4 for the definition of the length of a measured lamination in a Kleinian surface manifold). But since $e^{-}$ and $\lambda^{+}$ fill $S(e^{-})$ and $\mu$ has support contained in $L\subset e^{-}$, we obtain that $i(\mu,\lambda^{+})>0$. We claim that this forces $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$ to be unbounded, a contradiction with Bridgeman’s result.
Since $i(\mu,\lambda^{+})>0$, there is a minimal component $\lambda^{+}_{0}\subset\lambda^{+}$ with $i(\mu,\lambda^{+}_{0})>0$. If $\lambda^{+}_{0}$ is a closed curve, then (4) implies that $(\ell_{\rho_{n}}(\lambda^{+}_{0}))_{n}$ is unbounded, so is $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$. Let us now assume that $\lambda^{+}_{0}$ is irrational. We cannot apply (4) to prove the claim since that implication is a priori only valid for closed curves (and, by linearity, for weighted multicurves). Better control on (4) is given by Otal’s continuity theorem [Ota96, Theorem 4.0.1], of which we need the more general form stated by Lecuire in [Lec20, Theorem 3.7].
@@
Let $(\rho_{n})_{n\in{\mathbb{N}}}\subset{\mathcal{AH}}(S)$ be a sequence converging to a small minimal action of $\pi_{1}(S)$ on a real tree ${\mathcal{T}}_{\mu}$, where $\mu\in{\mathcal{ML}}(S)$. Let $\epsilon_{n}\searrow 0$ be such that $\epsilon\ell_{\rho_{n}}\to i(\mu,-)$. Let $s_{0}\in{\mathcal{T}}(S)$ be any hyperbolic metric on $S$. Let $\alpha\subset S$ be a geodesic lamination realized in ${\mathcal{T}}_{\mu}$, i.e. each leaf of $\alpha$ intersects $\mu$ transversely. Then there exists a neighborhood ${\mathcal{V}}(\alpha)\subset{\mathcal{GL}}(S)$ of $\alpha$ and constants $K$, $n_{0}$ such that, for any simple closed curve $c\subset{\mathcal{V}}(\mu)$ and for any $n\geq n_{0}$,
@@
Pick a sequence of weighted simple closed curves $(\alpha_{j})_{j\in{\mathbb{N}}}$ converging to $\lambda^{+}_{0}$ in ${\mathcal{ML}}(S)$. By Otal’s continuity theorem there exist constants $K$, $N$ and $J$ such that, for all $n\geq N$ and $j\geq J$,
$\ell_{\rho_{n}}(\alpha_{j})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\alpha_{j}),$
where $s_{0}$ is any fixed reference metric, independent of $n$ and $j$. By continuity of length functions (Theorem 2.6 by Brock [Bro00]), the limit $j\to\infty$ gives, for each $n\geq N$,
$\ell_{\rho_{n}}(\lambda^{+}_{0})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\lambda^{+}_{0}).$
Since $\epsilon_{n}\searrow 0$, while $K\ell_{s_{0}}(\lambda^{+}_{0})$ is a positive constant, we obtain that $\ell_{\rho_{n}}(\lambda^{+}_{0})\to\infty$, concluding the proof. ∎

### v2

@@ 1. Introduction
[…]
Figure 1. Schematic representation of hyperbolic manifolds homeomorphic to $S\times{\mathbb{R}}$, with 11-dimensional “$S$-direction”, conformal boundary in red and convex core in grey. The structures are, from left to right: Fuchsian, quasi-Fuchsian, Kleinian with a cusp and Kleinian with simply-degenerated bottom end.
@@
More precisely, let $P\subset S^{-}$ be a simple multicurve, let $\Sigma^{-}$ be a union of connected components of $S^{-}-P$ and let $L$ be a union of arational laminations on each components of $S^{-}-(P\cup\Sigma^{-})$. Denote by ${\mathcal{AH}}_{o}^{+}(S;P\cup L)\subset{\mathcal{AH}}_{o}(S)$ the subspace of one-sided degenerated structures with bottom end cusped along $P$ and having ending laminations $L$. Let ${\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})\subset{\mathcal{ML}}(S^{+})$ be the subspace of measured laminations $\mu$ on $S^{+}$ without closed leaves of weight $\geq\pi$ and such that $\mu$ and $P\cup L$ fill $S-\Sigma^{-}$, i.e. any essential closed curve on $S$ intersects either $\mu\cup P\cup L$ or $\Sigma^{-}$.
@@ 1.1. Outline of the proof
First, we prove continuity in Section 3, see Proposition 3.4. Then, in section 4 we prove properness. The argument follows closely the cases considered by Bonahon-Otal [BO04], Lecuire [Lec06], Anderson-Lecuire [AL13], Baba-Ohshika [BO23] and Lecuire [Lec25]. The first step is to prove that any sequence of representations with convergent image under $\operatorname{m}^{-}\times\operatorname{b}^{+}$ must have an algebraically convergent subsequence, using the theory of degenerescence of metrics towards actions on real trees of Morgan-Shalen [MS84, Ota96]. Then, it follows from results of Anderson-Lecuire [AL13] and Bonahon-Otal and Lecuire [BO04, Lec06] that this algebraic limit must lie in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
@@
Consider a ray $r\colon[0,\infty)\to{\mathcal{T}}(S^{-})$ such that $\lim_{t\to\infty}r(t)={\mathcal{E}}^{-}$ in the space of end structures ${\mathcal{ES}}(S^{-})$ (called space of gallimaufries in [AL13], see Section 2.8 for a description of its topology), in other words, $r$ extends to a path $\overline{r}\colon[0,\infty]\to{\mathcal{ES}}(S^{-})$ with $\overline{r}(\infty)={\mathcal{E}}^{-}$. The objects and maps involved in the proof are summarized in the diagram below, where $\overline{\operatorname{U}}$ denotes the extension of $\operatorname{U}$ to the space of end structures.
$0,\infty)\times{\mathcal{T}}(S^{+})$${\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})$${\mathcal{QF}}_{o}(S)$${\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})$$0,\infty]\times{\mathcal{T}}(S^{+})$${\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})$${\mathcal{AH}}_{o}^{+}(S)$${\mathcal{ML}}(S^{+})$$\scriptstyle{r\times\rm{incl}}$$\scriptstyle{\operatorname{U}}$$\scriptstyle{\operatorname{b}^{+}\times\operatorname{b}^{-}}$$\scriptstyle{\rm{proj}_{2}}$$\scriptstyle{\overline{r}\times\rm{incl}}$$\scriptstyle{\operatorname{B}^{+}}$$\scriptstyle{\overline{\operatorname{U}}}$$\scriptstyle{\operatorname{b}^{+}}$
@@ 1.2. Further directions
The essential difficulty lies in proving that any geometrically finite structure on $S\times{\mathbb{R}}$ is uniquely determined by its pair of bending laminations. This is expected to hold, and it was proved in the case of rational bending laminations by Bonahon and Otal [BO04]. But it is not yet known in the general case, where the current strategy needs to be refined.
@@ 2.2. Laminations and trees
A transverse measure on a geodesic lamination $L\in{\mathcal{GL}}(S)$ is a map $\mu$ associating to each closed arc transverse to $L$ a number $\mu(k)\in{\mathbb{R}}_{\geq 0}$ (also denoted $i(\mu,k)$) which is invariant under transverse isotopy of arcs, countably additive, and with full support, i.e. $\mu(k)>0$ if and only if $k\cap L\not$$\emptyset$. Let ${\mathcal{ML}}(S)$ denote the space of measured lamination on $S$, endowed with the weak topology. The geodesic laminations that admit a transverse measure (with full support) are precisely the recurrent laminations [PH92, §1.7]. A minimal geodesic lamination admits a unique transverse measure up to scaling. See [PH92] for more details on measured laminations, and a description of a piecewise linear structure on ${\mathcal{ML}}(S)$ using train tracks. By Thurston earthquake theorem [Thu86], the space ${\mathcal{ML}}(S)$ is homeomorphic to ${\mathcal{T}}(S)$. See also [CEM06], [Ota96, Appendix], [MT98, §6.1].
@@ 2.4. Pleated surfaces
Let $M$ be a hyperbolic three-manifold, $S$ a complete hyperbolic surface (not necessarily compact) and $L$ a geodesic lamination on $S$. A pleated surface with pleating locus $L$ is a continuous map $f\colon S\to M$ which is a totally geodesic immersion on SLS-L and on $L$. In general, the flat pieces of SLS-L are bent along the pleating locus, with an angle recorded into a bending cocycle $\beta$ which is an element of the space ${\mathcal{H}}(L;{\mathbb{R}}/2\pi{\mathbb{Z}})$ of ${\mathbb{R}}/2\pi{\mathbb{Z}}$-valued cocycles along $L$ [Bon96] (which is a torus whose dimension is determined by $S$ and $L$). When $f$ is locally convex, this bending cocycle defines a transverse measure with support contained in $L$, called the bending or pleating (measured) lamination of $f$. The induced path metric on $S$ inside $M$ is hyperbolic and defines a point $h\in{\mathcal{T}}(S)$, which can also be encoded in a shearing cocycle $\sigma\in{\mathcal{H}}(L;{\mathbb{R}})$. The boundary of the convex core $C(\rho)$ of $\rho\in{\mathcal{AH}}(S)$ is an embedded convex pleated surface [CEM06, II.1.11,12]. In particular, its bending measured lamination has weight $\pi$ along closed leaves.
@@ 2.7. End structures
Baba and Ohshika described the combination of ending laminations, parabolic locus and bending laminations that occur for Kleinian surface groups [BO23]. It follows from their result that, for $\rho\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ where $e^{-}=P\cup L$, the bending lamination $\operatorname{b}^{+}(\rho)$ of the top boundary of its convex core fills $S(e^{-})$ with $e^{-}$, i.e. any essential closed curve on $S(e^{-})$ intersects $P\cup L$ or $\operatorname{b}^{+}(\rho)$ (in other words, any essential closed curve on $S$ intersects $P\cup L\cup\Sigma^{-}\cup\operatorname{b}^{+}(\rho)$). Hence it lies in the subspace ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ consisting of those measured laminations without closed leaves of weight $\geq\pi$, and which fill $S(e^{-})$ with $e^{-}$.
@@ 2.8. Metrics converging to an end structure
In [AL13], Anderson and Lecuire define a natural topology on the space ${\mathcal{ES}}(S)$ of end structures on $S$, i.e. of doubly incompressible quadruplets $(P,\Sigma,m,L)$ on $S$ as in the previous section. For our purpose we only need to describe one particular type of convergence. A sequence of hyperbolic metrics $m_{n}\in{\mathcal{T}}(S)$ (seen as end structures of the form $(\emptyset,\emptyset,m_{n},\emptyset)$) converges to an end structure ${\mathcal{E}}=(P,\Sigma,m,L)$ if the following two conditions hold :
@@
Thus, we can identify ${\mathcal{T}}(S)$ with an open subspace of ${\mathcal{ES}}(S)$. Note that ${\mathcal{ES}}(S^{-}\sqcup S^{+})={\mathcal{ES}}(S^{-})\times{\mathcal{ES}}(S^{+})$.
The end invariants of an element $\rho$ of ${\mathcal{AH}}_{o}(S)$ form an end structure $\operatorname{E}(\rho)$ in ${\mathcal{ES}}(S^{-}\sqcup S^{+})$. For $\rho\in{\mathcal{AH}}_{o}^{+}(S)$, its end structure $\operatorname{E}(\rho)$ has the form $(P,\Sigma^{-},m^{-},L)\times(\emptyset,\emptyset,m^{+},\emptyset)\in{\mathcal{E}}(S^{-})\times{\mathcal{T}}(S^{+})$ (and any such end structure is doubly incompressible in the sense of [AL13, §2.9]).
@@ 4. Properness
The goal of this section is to prove that the map $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper (see the end of Section 2.7 for a description of the bending laminations that occur for strucutes in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$). This will imply that $\operatorname{b}^{+}_{\infty}\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper too. The proof is very similar to the proof of properness of similar maps by Baba and Ohshika in [BO23], Anderson and Lecuire in [AL13] and Lecuire in [Lec25]. It essentially follows by combining the last two references.
@@
Let ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ be the subspace of ${\mathcal{ML}}(S^{+})$ consisting of the measured laminations $\mu$ satisfying the following conditions :
@@ 4.2. Algebraic convergence
[…]
@@
Since $e^{-}$ and $\lambda^{+}$ fill $S(e^{-})$ and $\mu$ has support contained in $L\subset e^{-}$, we obtain that $i(\mu,\lambda^{+})>0$. Thus there is a minimal component $\lambda^{+}_{0}\subset\lambda^{+}$ with $i(\mu,\lambda^{+}_{0})>0$. In other words, the support of $\lambda^{+}_{0}$ is realized in ${\mathcal{T}}_{\mu}$.
This leads to a contradiction, as Lecuire shows in [Lec25]. His argument is divided into two cases, depending on whether $\lambda^{+}_{0}$ is supported on an irrational lamination [Lec25, §3.1] or a closed curve [Lec25, §3.2]. The resulting contradictions stem from Otal’s continuity theorem applied to $\lambda^{+}_{0}$ [Ota96, Theorem 4.0.1][Lec20, Theorem 3.7], which provides uniform control on the growth of lengths of curves (or weighted multicurves, by linearity) that are close to a lamination realized in the limit tree. Using convergence of the bending laminations $(\lambda^{+}_{n})_{n}$ to $\lambda^{+}$, Lecuire constructs sequences of weighted simple closed curves whose lengths with respect to $\rho_{n}$ grow too slowly, leading to a contradiction.
@@
Let $(\rho_{n})_{n\in{\mathbb{N}}}\subset{\mathcal{AH}}(S)$ be a sequence converging to a small minimal action of $\pi_{1}(S)$ on a real tree ${\mathcal{T}}_{\mu}$, where $\mu\in{\mathcal{ML}}(S)$. Let $\epsilon_{n}\searrow 0$ be such that $\epsilon\ell_{\rho_{n}}\to i(\mu,-)$. Let $s_{0}\in{\mathcal{T}}(S)$ be any hyperbolic metric on $S$. Let $\alpha\subset S$ be a geodesic lamination realized in ${\mathcal{T}}_{\mu}$, i.e. each leaf of $\alpha$ intersects $\mu$ transversely. Then there exists a neighborhood ${\mathcal{V}}(\alpha)\subset{\mathcal{GL}}(S)$ of $\alpha$ and constants $K>0$, $n_{0}$ such that, for any simple closed curve $c\in{\mathcal{V}}(\alpha)$ and for any $n\geq n_{0}$,
@@
[…]
In the second case, $\lambda_{0}^{+}$ is supported on a simple closed curve $d$. In particular, $\epsilon_{n}\ell_{\rho_{n}}(d)\to i(\mu,d)>0$. In that case, Lecuire shows that any closed curve $c$ transverse to $d$ satisfies [Lec25, Lemma 3.5]
$\frac{\epsilon_{n}\ell_{\rho_{n}}(d)}{\epsilon_{n}\ell_{\rho_{n}}(c)}=\frac{\ell_{\rho_{n}}(d)}{\ell_{\rho_{n}}(c)}\to 0.$
But $\epsilon_{n}\ell_{\rho_{n}}(c)\to i(\mu,c)$, so $\epsilon_{n}\ell_{\rho_{n}}(d)\to 0$, a contradiction. This concludes the proof. ∎